The scale drawing is 4 cm:5 m. Give the length, width, and area of the actual room. ( 8 cm & 12 cm

1
Expert's answer
2022-04-07T14:41:56-0400

Length: "L=\frac{12}{4}*5=15m."

Width: "W=\frac{8}{4}*5=10m."

Area: "A=15*10=150m."